OpenStudy (albert0898):

30. If 8 sin^2 θ + 2 sin θ − 1 = 0, then what is the smallest positive value of θ? (A) 7.3° (B) 14.5° (C) 30° (D) 60° (E) 75.5° What are the steps to get the answer?

Nnesha (nnesha):

that's same as quadratic equation Ax^2+Bx+C=0 let sin theta =y you can rewrite it as \[\huge\rm 8y^2+ 2 y − 1 = 0\] can you find factor of this equation ?

OpenStudy (albert0898):

y^2 + 2y - 8 = 0 (y+4)(y-2)=0 y = -4, 2

OpenStudy (albert0898):

So, how do I figure out the degrees?

Nnesha (nnesha):

hmm you can't just move the leading to constant term ...

Nnesha (nnesha):

you need to multiply AC but that doesn't mean you can write 8 as the constant term.

Nnesha (nnesha):

AC= leading coefficient times constant term so 8 times -1=-8 now find two numbers whose product equal to AC and sum should be equal to coefficient of y term

OpenStudy (albert0898):

So AC = -8 Sum = 2 +4, -2

Nnesha (nnesha):

correct now since the leading coefficient isn't one try group by factoring method  keep the first and constant term and write both numbers in the middle by factor by grouping